Write an equation of the line, in slope-intercept form, that passes through the points (0,4) and (-1,3)
Answer by subudear(62)

Line equation is given by y = mx +c
m = (y2 - y1)/(x2 - x1)
lets say (x1, y1) = (0,4) and (x2,y2) = (-1,3)
then m = (3 - 4)/(-1 - 0) = 1
equation of line is
y - 4 = (1)(x -0)
y - 4 = x
y = x + 4